OpenStudy (anonymous):

If the function f is continuous for all real numbers and f(x) = (x^2-4)/(x+2) when x cannot equal -2 , then f(-2) =? ....I would have plugged in the f but it said x cannot equal to -2 so Im confused.

OpenStudy (anonymous):

you could always factor your numerator ;-) but that would kinda defeat the purpose I'm guessing. Perhaps using l'hopital's rule here? if you replace x by -2, you get 0/0

OpenStudy (anonymous):

Yeah, 0 was the only thing I could come up with but my calculus teacher always claims I do illegal math so I was making sure I wasnt making up something.

OpenStudy (anonymous):

I'm under the impression that you seem to be thinking that the answer is 0. Because 0/0 is NOT equal to 0 (well, sometimes, it might, but it's an exception). 0/0 is an undetermined form, which you need to apply l'Hopital's rule to determine the value of

OpenStudy (anonymous):

Easiest thing you can do is factorize really : x^2-4 = (x+2)(x-2) so f(x) = (x+2)(x-2)/(x+2) = x-2. f(-2) = -2 -2 = -4

OpenStudy (anonymous):

So I used the defference of squares to factor the top and I understand the (x=2)(x-2)(x=2) but I dont understand how you made that equal to (x+2). I divided by (x+2/0 instead.

OpenStudy (anonymous):

(x+2)*

OpenStudy (anonymous):